40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 238     Entry time: Wed Jul 28 19:46:46 2010
Author: Megan 
Type: DailyProgress 
Category:  
Subject: PSL Progress! 

The PMC is still locked (except when we do something to unlock it) and both the analyzer cavities and reference cavities have the ability to be locked. We have not locked them simultaneously yet but will be working on that so we can get a beat. We connected the VCO driver to the AOM for the analyzer cavity and got it locked.

A preliminary calibration for the photodetectors in the path of the transmitted light from the reference cavity has been calculated, although changes in the next day or two might change it, so it has not been put into the computer.

The table by the cavities is now very pretty . Until we fill all the space we just cleared with everything we need to beat the two beams together!

Preparations for that will be started tomorrow by putting the correct optics in the path of the transmitted light of the analyzer cavity and better aligning those in the path of the reference cavity's transmitted light.

Also, as an extra note, last week we calibrated some inputs, so we no longer have anti-energy reflected from the PMC! In fact, at this moment, we have 1.4mW. Success is still 100 and strength is down to 44! Guess it's time for dinner.

ELOG V3.1.3-